A 1.55 g

1.55 g sample of an unknown gas at 67 

C

67 ∘C and 1.05 atm

1.05 atm is stored in a 2.95 L

2.95 L flask.

What is the density of the gas?


1
Expert's answer
2022-04-05T10:04:01-0400

PV = nRT

P = 1.05atm

V = 2.95L

R = 0.08206Latm/Kmol

T = 67 + 273 = 340K

n = moles = PV/RT

Use PV = nRT to determine moles of gas present in gas:

(1.05 atm) (2.95 L) = (n) (0.08206) (340 K)

n = 0.111 mol

2) Get molar mass of gas using calculated moles:

0.111 mol x M g/mol = 1.55g

M = 1.55 g / 0.111 mol = 13.96 g/mol

Density = Molar Mass x P / RT

= 13.96 g/mol x 1.05 atm / 0.08206 L mol / atm .K x 340 K

= 0.525 g /L